OpenStudy (anonymous):

The value of Jamie's baseball card collection rose 50% in the last year. It is now worth $255. What was it worth a year ago?

OpenStudy (kropot72):

$255 is 150% of the value a year ago. 1% of the value a year ago = 255/150 100% of the value a year ago = (255/100) * 100 = you can calculate